An evolutionary biologist once said: “Would I lay down my life to save my brother? No, but I would to save two brothers or eight cousins.” Explain what he meant by this. - Why does this 'make sense'.
Answer #1

Ans: For explanning this statement we have to go through coefficient of relatedness (r) which is the probability that two individuals share an allele due to recent common ancestry.

In human relationships, the value of the coefficient of relationship is usually calculated based on the knowledge of a full family tree extending to a comparatively small number of generations, perhaps of the order of three or four.

r( coefficient of relatedness)= 2-n; n= degree of relationship.

Degree of
relationship
Relationship Coefficient of
relationship (r)
0 identical twins; clones 100% or 1
1 mother / father / daughter / son 50% (2−1)
1
parent's identical twin / identical twin's child
50% (2−1)
2 full sister / full brother 50% (2⋅2−2)
2 grandmother / grandfather / granddaughter / grandson 25% (2−2)
4 Cousin 12.5% (2-4)

So genetically speaking, two brothers (summation of r= 2*0.5= 1) is the same as eight cousins (summation of r=8*0.125=1).

So he meant that in context of coefficient of relationship 2 brothers are equivalent to 8 cousins.